OpenStudy (anonymous):

above zero degrees, the actula number of degrees celcius for any given temperature is _______ the degrees farenheit

OpenStudy (anonymous):

Lower than 0 degrees Celsius is equal to 32 degrees Fahrenheit

OpenStudy (anonymous):

it gives me mulitple choice answers it says a. higher than b. lower than c. the same as or d. a or b i am a little confuesd

OpenStudy (anonymous):

it gives me mulitple choice answers it says a. higher than b. lower than c. the same as or d. a or b i am a little confuesd

OpenStudy (anonymous):

Above zero degrees, the temperature in Celsius is LOWER than that of the corresponding Fahrenheit (Answer is B) --> the reason has to do with the formula. (°C × 9/5) + 32 = °F So: Fahrenheit will always be a larger value than the Celsius (even at zero degrees F = C+32). eg. 0° C = 32° F 500° C = 932°F